# Individual or collective treatments: how to target antimicrobial use to limit the spread of Mannheimia haemolytica among beef cattle?

### v2

@@ Results
Figure 2: Distributions of the cumulative incidence of BRD cases in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
@@
Collective treatments successfully reduced the mortality in HR90 scenarios when compared to individual treatments (Figure 5). The effect was however more contrasted in HR30 scenarios, with only the mortality associated with the conventional collective treatment being significantly lower than the individual treatment in both Sorted and Random scenarios (p-value <0.05). The statistical comparisons within collective treatment strategies yielded unconclusive results with regards to significance.
Figure 5: Boxplots of the mortality in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
@@ Discussion
This study aimed at modeling the impact of triggering collective antimicrobial treatments according to three different criteria on the spread of M. haemolytica in multi-pens fattening systems while accounting for the multi-pen structure, individual risk of being affected by BRD and farming practices in terms of pens composition. We computed the cumulative incidence, the average duration of severe clinical signs and the mortality. We put these outputs in balance with the AMU and treatment precision for each scenario. We also computed the delay between the beginning of an episode and the beginning of the treatment and calculated the proportion of pens triggering collective treatments based on each criterion. We proposed that a novel criterion, based on the speed of the spread rather than on a proportion of detected individuals, could yield a better trade-off between BRD case incidence and AMU than the conventional way of implementing collective treatment. This result could have interesting operational implications on the decision-making regarding collective treatment in fattening farms.
@@
We showed that collective treatment could efficiently reduce the incidence of bacterial respiratory disease in population with high individual risk of BRD. The positive impact on animal health and welfare balanced the supplementary cost in antimicrobial doses. This compromise in favour of reducing BRD impact was particularly relevant and valuable for high-risk individuals. This finding is consistent with a previous simulation study which implemented collective treatments on an âaverage pathogenâ with less strict thresholds ([picault_modelling_2022]). In American systems, medication of at-risk cattle is frequent at feedlot entry after their idenfication and the implementation of preconditionning programs ([gonzalez-martin_reducing_2011]). These programs has been studied in Europe, but the results remained inconclusive regarding their positive impact on animal health and welfare ([vanbergue_comparison_2024]). The study however concluded that the cattle sector should focus its effort on bettering husbandry practices. This suggestion aligns with the use of a selective criterion for the implementation of curative collective treatments.
@@
Our results showcased a positive impact of collective treatments on disease related mortality in the herd in high risk scenarios. However, no criterion was identified as systematically better than the others in reducing the mortality. This can be due to the fact that collective treatments happened rather rapidly after an outbreak, in spite of some differences in delay. This swift reaction did not let time for animals to develop severe clinical signs and to die subsequently to them. Our results aligned with field studies showingreduced mortality in herds where collective treatment was used ([oconnor_mixed_2013]).
@@
The supplementary cost in antimicrobial was less desirable when the general risk was lower in the population, as the reduction in incidence and severity thanks to collective treatment was less substantial. The overuse of antimicrobial leads to antimicrobial resistance and treatment purchase and administration are costs supported by the farmer ([laxminarayan_antibiotic_2013]). Therefore, there appears to be a trade-off between incidence reduction and antimicrobial use which could benefit from a characterization of the risk level of the pens. Such a characterization could guide the farmers in making the best decision regarding treatment and herd management ([edwards_control_2010]).
@@
Our model allowed us to identify a beneficial novel criterion for collective treatment. Indeed, instead of a threshold of detected animals in a pen, basing a collective treatment decision on the speed of the spread of the disease allowed marginal gains in disease incidence while drastically reducing the AMU and the treatment misuse when compared to the other collective treatment criteria. A field meta-analysis of randomized clinical trials showed that the risk reduction of BRD varied drastically with the criterion for collective treatment ([baptiste_antimicrobial_2017]). The routine criterion is 10% of morbidity for two consecutive days, however, there is no evidence based studies proving the relevance of this threshold ([radostits_herd_2001]). Perspectives of this study include the investigation of the novel treatment strategies in broader contexts in order to translate the thresholds used into operative indicators.
@@
Collective treatments based on the speed of the spread of the pathogen proved to be triggered sooner than with the other criteria. This precocity was key for the reduction of the incidence and severity of the disease, especially in high-risk populations. It also showed to target more specifically high-risk groups than the other criteria, as it reduced the proportion of low-risk pens having received collective treatments. On an individual level, the delay between the infection and the administration of the first dose of treatment was also characterized. This delay was shortened by the implementation of collective treatment strategies, particularly for the conventional criterion. These are original findings with clear practical implications, especially with regards to the need for a better knowledge by farmers of individual risk of animals developing BRD, in order to account for this information when constituting pens. However, triggering collective treatments greatly relies on the ability to detect diseased animals in the first place, therefore, increasing the sensitivity via standardized detection procedures may increase the benefit of the subsequent control measures and may modify the most convenient criteria to trigger collective treatment ([kamel2024strategies]).
@@
Our model simulated scenarios assuming the circulation of only one bacterial pathogen. BRD is essentially a multi-pathogen disease, and synergies between viruses and bacteria have been reported to change the disease dynamics and severity ([gaudino_understanding_2022]) . Further investigations of the effect of treatment strategies could be conducted in a context with mixed infections. However, quantifying the treatment response in such context is still a pending question ([fulton_lung_2009]).
There is clear evidence that antimicrobial resistance is rising among M.haemolytica populations ([woolums_multidrug_2018, timsit_prevalence_2017]). This rise happens on time scales that are longer than the one we modeled. Therefore, we modeled the existence of resistant strains by calculating a probability of treatment success, which value was set along the span of the simulation. Future work could aim at evaluating the long term effects of collective treatment protocols on the antimicrobial trends.
The force of infection used in this model assumed an increased pathogen transmission within pens with an additional equal contribution of the surrounding pens. This hypothesesis relied on the fact that aerosolization of M.haemolytica has been pointed as a contributing factor to rapid disease spread among cattle populations in close quarters ([rice_mannheimia_2007]). However, this airborne transmission is not as effective as transmission via close contacts. Adaptions of the model could therefore implement a force of infection based on the farm spatial organization ([ackermann2000response]).
@@ Acknowledgements
We thank our colleagues from the BIOEPAR Dynamo team for their comments and advice on this work. We are most grateful to the INRAE MIGALE bioinformatics facility ([noauthor_migale_nodate]), for the use of their computing resources.
@@
The complete code for the model model can be found on this git repository: https://archive.softwareheritage.org/browse/origin/directory/?branch=refs/heads/PCI_animal_science2024&origin_url=https://forgemia.inra.fr/dynamo/brd-models/brd-public.git. The version of the model used for this work is tagged under pci_animal_science_2024.
@@ B.1 Method
In order to determine the relevant threshold for intervention, we defined a range of viable threshold values. We thus plotted both criteria on baseline scenarios (without any collective treatments). This baseline scenarios was repeated along 200 stochastic replicates in order to obtain a distribution of values for each time step. For each time step, the median value, the 5th and 95th percentiles were recorded, thus resulting in three time series for each criterion and each scenario (Figures 13,14).
@@
The interval was extracted from the HR30 Random scenario for both criteria. The upper bound was defined as the maximum point of the 95% quantile time series and the lower bound was defined by the maximum point of the median time series. The rationale behind this was to exclude values that would either be too low or too likely while also excluding any values that would be too high or too unlikely. Indeed, HR30 Random represents a baseline scenario and our selection allowed us to select only the higher half of the maximum values in this scenario. Thus, we obtained a credible range of [1,3]delimited-[]1.3[1,3][ 1,3 ] for the severity criterion, discretized with intervals of 0.2. This criterion is also characterized by parameter $\alpha$, assumed strictly lower than 0.5 to emphasize the number of individuals with severe clinical signs. This parameter was assigned a credible range of [0.1,0.4]delimited-[]01.04[0.1,0.4][ 0.1,0.4 ] with intervals of 0.1.
We used the same rationale for the selection of a credible range for the slope criterion. We thus selected a range encompassing the maxium value of the median run in the HR30-Random scenario and the maxium of the 95% quantile of this scenario. For the slope criterion, the credible range was thus [0.75,1.7]delimited-[]075.17[0.75,1.7][ 0.75,1.7 ], discretized by intervals of 0.05.

### v4

@@ Results
Figure 2: Distributions of the cumulative incidence of BRD cases in the building population. The boxplots feature the median and quartiles of the distributions. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals.
@@
Collective treatments successfully reduced the occurence of high mortality in HR90 scenarios when compared to individual treatments (Figure 5). Indeed, in random scenarios, the mortality was higher or equal to é% in around 28% or scenarios with individual treatment, whereas this mortality was obtained in only 10 to 13% of replicates with collective treatments. In HR30 scenarios, the differences of in high mortality occurence are less marked, as most of the replicates yielded a mortality lower or equal to 1%, regardless of the applied treatment strategy.
@@
Figure 5: Barplots of the mortality in the building population. First column: the animals are allocated randomly in pens. Second column: the animals are allocated in pens according to their individual risk level. Top row : 30, 40 and 30% of low, medium and high risk animals respectively. Bottom row: 10% of low risk animals and 90% of high risk animals
@@ Discussion
This study aimed at modeling the impact of triggering collective antimicrobial treatments according to three different criteria on the spread of M. haemolytica in multi-pens fattening systems while accounting for the multi-pen structure, individual risk of being affected by BRD and farming practices in terms of pens composition. We computed the cumulative incidence, the average duration of severe clinical signs and the mortality. We balanced these outputs against the AMU and treatment precision for each scenario. We also computed the delay between the beginning of an episode and the beginning of the treatment and calculated the proportion of pens triggering collective treatments based on each criterion. We proposed that a novel criterion, based on the speed of the spread rather than on a proportion of detected individuals, could yield a better trade-off between BRD case incidence and AMU than the conventional way of implementing collective treatment. This result could have interesting operational implications on the decision-making regarding collective treatment in fattening farms.
@@
We showed that collective treatment could efficiently reduce the incidence of bacterial respiratory disease in population with high individual risk of BRD. The positive impact on animal health and welfare balanced the supplementary cost in antimicrobial doses. This compromise in favour of reducing BRD impact was particularly relevant and valuable for high-risk individuals. This finding is consistent with a previous simulation study which implemented collective treatments on an âaverage pathogenâ with less strict thresholds ([38]). In American systems, medication of at-risk cattle is frequent at feedlot entry after their idenfication and the implementation of preconditionning programs ([20]). These programs has been studied in Europe, but the results remained inconclusive regarding their positive impact on animal health and welfare ([50]). The study however concluded that the cattle sector should focus its effort on improving husbandry practices. This suggestion aligns with the use of a selective criterion for the implementation of curative collective treatments.
@@
Our results showcased a positive impact of collective treatments on disease related mortality in the herd in high risk scenarios. However, no criterion was identified as systematically better than the others in reducing the mortality. This can be due to the fact that collective treatments happened rather rapidly after an outbreak, in spite of some differences in delay. This swift reaction did not let time for animals to develop severe clinical signs and to die subsequently to them. Our results aligned with field studies showing reduced mortality in herds where collective treatment was used ([36]).
@@
The supplementary cost in antimicrobial was less desirable when the general risk was lower in the population, as the reduction in incidence and severity thanks to collective treatment was less substantial. The overuse of antimicrobial leads to antimicrobial resistance and treatment purchase and administration are costs supported by the farmer ([30]). Therefore, there appears to be a trade-off between incidence reduction and antimicrobial use which could benefit from a characterization of the risk level of the pens. Such a characterization could guide the farmers in making the best decision regarding treatment and herd management ([15]).
@@
Our model allowed us to identify a beneficial novel criterion for collective treatment. Indeed, instead of a threshold of detected animals in a pen, basing a collective treatment decision on the speed of the spread of the disease allowed marginal gains in disease incidence while drastically reducing the AMU and the treatment misuse when compared to the other collective treatment criteria. A field meta-analysis of randomized clinical trials showed that the risk reduction of BRD varied drastically with the criterion for collective treatment ([6]). The routine criterion is 10% of morbidity for two consecutive days, however, there is no evidence based studies proving the relevance of this threshold ([40]). Perspectives of this study include the investigation of the novel treatment strategies in broader contexts in order to translate the thresholds used into operative indicators.
@@
Collective treatments based on the speed of the spread of the pathogen proved to be triggered sooner than with the other criteria. This precocity was key for the reduction of the incidence and severity of the disease, especially in high-risk populations. It also showed to target more specifically high-risk groups than the other criteria, as it reduced the proportion of low-risk pens having received collective treatments. On an individual level, the delay between the infection and the administration of the first dose of treatment was also characterized. This delay was shortened by the implementation of collective treatment strategies, particularly for the conventional criterion. These are original findings with clear practical implications, especially with regards to the need for a better knowledge by farmers of individual risk of animals developing BRD, in order to account for this information when constituting pens. However, triggering collective treatments greatly relies on the ability to detect diseased animals in the first place, therefore, increasing the sensitivity via standardized detection procedures may increase the benefit of the subsequent control measures and may modify the most convenient criteria to trigger collective treatment ([26]).
@@
Our model simulated scenarios assuming the circulation of only one bacterial pathogen. BRD is essentially a multi-pathogen disease, and synergies between viruses and bacteria have been reported to change the disease dynamics and severity ([19]) . Further investigations of the effect of treatment strategies could be conducted in a context with mixed infections. However, quantifying the treatment response in such context is still a pending question ([18]).
There is clear evidence that antimicrobial resistance is rising among M.haemolytica populations ([53, 48]). This rise happens on time scales that are longer than the one we modeled. Therefore, we modeled the existence of resistant strains by calculating a probability of treatment success, which value was set along the span of the simulation. Future work could aim at evaluating the long term effects of collective treatment protocols on the antimicrobial trends.
The force of infection used in this model assumed an increased pathogen transmission within pens with an additional equal contribution of the surrounding pens. This hypothesesis relied on the fact that aerosolization of M.haemolytica has been pointed as a contributing factor to rapid disease spread among cattle populations in close quarters ([41]). However, this airborne transmission is not as effective as transmission via close contacts. Adaptions of the model could therefore implement a force of infection based on the farm spatial organization ([2]).
@@ Acknowledgements
We thank our colleagues from the BIOEPAR Dynamo team for their comments and advice on this work. We are most grateful to the INRAE MIGALE bioinformatics facility ([32]), for the use of their computing resources.
@@
The complete code for the model model can be found on this git repository: https://forgemia.inra.fr/dynamo/brd-models/brd-public/-/tree/PCI_animal_science2024. The version of the model used for this work is tagged under pci_animal_science_2024.
@@ B.1 Method
In order to determine the relevant threshold for intervention, we defined a range of viable threshold values. We thus plotted both criteria on baseline scenarios (without any collective treatments). This baseline scenarios was repeated along 200 stochastic replicates in order to obtain a distribution of values for each time step. For each time step, the median value, the 5th and 95th percentiles were recorded, thus resulting in three time series for each criterion and each scenario (Figures 12,13).
@@
The interval was extracted from the HR30 Random scenario for both criteria. The upper bound was defined as the maximum point of the 95% quantile time series and the lower bound was defined by the maximum point of the median time series. The rationale behind this was to exclude values that would either be too low or too likely while also excluding any values that would be too high or too unlikely. Indeed, HR30 Random represents a baseline scenario and our selection allowed us to select only the higher half of the maximum values in this scenario. Thus, we obtained a credible range of $[1,3]$ for the severity criterion, discretized with intervals of 0.2. This criterion is also characterized by parameter $\alpha$, assumed strictly lower than 0.5 to emphasize the number of individuals with severe clinical signs. This parameter was assigned a credible range of $[0.1,0.4]$ with intervals of 0.1.
We used the same rationale for the selection of a credible range for the slope criterion. We thus selected a range encompassing the maxium value of the median run in the HR30-Random scenario and the maxium of the 95% quantile of this scenario. For the slope criterion, the credible range was thus $[0.75,1.7]$, discretized by intervals of 0.05.
@@ B.2.2 Slope criterion
Annexe C Additional figures
Figure 19: Impact of the collective treatment strategies on the outputs with the tested thresholds. Top panel: Cumulative incidence. Bottom panel: AMU. First row: the animals are allocated randomly in pens. Second row: the animals are allocated in pens according to their individual risk level. Each column represents the tested value for the transmission rate. Statistical test: Wilcoxonn-Mann-Whitney test with Bonferroni’s correction.
1
  • eng 1
    • Elsevier 1
      • en 1
        • en 1
          • Wiley Online Library 1
            • English 1
              • en 1
                • en 1
                  • en 1
                    • en 1
                      • eng 1
                        • en 1
                          • en 1
                            • en 1
                              • en 1
                                • eng 1
                                  • MDPI 1
                                    • en 1
                                      • en 1
                                        • eng 1
                                          • en 1
                                            • eng 1
                                              • eng 1
                                                • Philadelphia 1
                                                  • Saunders Philadelphia 1
                                                    • eng 1
                                                      • Elsevier 1
                                                        • en 1
                                                          • en 1
                                                            • American Association for the Advancement of Science 1
                                                              • en 1
                                                                • eng
                                                                  Références
                                                                  • [1] K.. Abell, M.. Theurer, R.. Larson, B.. White and M. Apley “A mixed treatment comparison meta-analysis of metaphylaxis treatments for bovine respiratory disease in beef cattle” In Journal of Animal Science 95.2, 2017, pp. 626–635 DOI: 10.2527/jas.2016.1062
                                                                  • [2] Mark R Ackermann and Kim A Brogden “Response of the ruminant respiratory tract to Mannheimia (Pasteurella) haemolytica” In Microbes and infection 2.9, 2000, pp. 1079–1088
                                                                  • [3] David E. Amrine, Jiena G. McLellan, Brad J. White, Robert L. Larson, David G. Renter and Mike Sanderson “Evaluation of three classification models to predict risk class of cattle cohorts developing bovine respiratory disease within the first 14â¯days on feed using on-arrival and/or pre-arrival information” In Computers and Electronics in Agriculture 156, 2019, pp. 439–446 DOI: 10.1016/j.compag.2018.11.035
                                                                  • [4] S. AssiÃľ, H. Seegers, B. Makoschey, L. DÃľsirÃľ-BousquiÃľ and N. Bareille “Exposure to pathogens and incidence of respiratory disease in young bulls on their arrival at fattening operations in France” Number: 7 In Veterinary Record 165.7, 2009, pp. 195–199 DOI: 10.1136/vr.165.7.195
                                                                  • [5] D Baggott, A Casartelli, F Fraisse, C Manavella, R Marteau, S Rehbein, M Wiedemann and S Yoon “Demonstration of the metaphylactic use of gamithromycin against bacterial pathogens associated with bovine respiratory disease in a multicentre farm trial” In Veterinary Record 168.9, 2011, pp. 241–241
                                                                  • [6] Keith Edward Baptiste and Niels Christian Kyvsgaard “Do antimicrobial mass medications work? A systematic review and meta-analysis of randomised clinical trials investigating antimicrobial prophylaxis or metaphylaxis against naturally occurring bovine respiratory disease” In Pathogens and Disease 75.7, 2017, pp. ftx083 DOI: 10.1093/femspd/ftx083
                                                                  • [7] Ken G. Bateman, S. Martin, Patricia E. Shewen and Paula I. Menzies “An evaluation of antimicrobial therapy for undifferentiated bovine respiratory disease” In The Canadian Veterinary Journal 31.10, 1990, pp. 689–696 URL: https://www.ncbi.nlm.nih.gov/pmc/articles/PMC1480842/
                                                                  • [8] Claudia Blakebrough-Hall, Paul Hick, T J Mahony and Luciano A GonzÃąlez “Factors associated with bovine respiratory disease case fatality in feedlot cattle” In Journal of Animal Science 100.1, 2022, pp. skab361 DOI: 10.1093/jas/skab361
                                                                  • [9] Calvin W. Booker, Sameeh M. Abutarbush, Paul S. Morley, G. Jim, Tom J. Pittman, Oliver C. Schunicht, Tye Perrett, Brian K. Wildman, R. Fenton, P. Guichon and Eugene D. Janzen “Microbiological and histopathological findings in cases of fatal bovine respiratory disease of feedlot cattle in western Canada” In The Canadian Veterinary Journal 49.5, 2008, pp. 473–481 URL: https://www.ncbi.nlm.nih.gov/pmc/articles/PMC2359492/
                                                                  • [10] Jeff L. Caswell, Ken G. Bateman, Hugh Y. Cai and Fernanda Castillo-Alcala “Mycoplasma bovis in Respiratory Disease of Feedlot Cattle” Number: 2 Publisher: Elsevier In Veterinary Clinics: Food Animal Practice 26.2, 2010, pp. 365–379 DOI: 10.1016/j.cvfa.2010.03.003
                                                                  • [11] Shih-Yu Chen, Pedro Negri Bernardino, Erik Fausak, Megan Van Noord and Gabriele Maier “Scoping Review on Risk Factors and Methods for the Prevention of Bovine Respiratory Disease Applicable to CowâCalf Operations” Number: 3 Publisher: Multidisciplinary Digital Publishing Institute In Animals 12.3, 2022, pp. 334 DOI: 10.3390/ani12030334
                                                                  • [12] Johann F Coetzee, Drew R Magstadt, Pritam K Sidhu, Lendie Follett, Adlai M Schuler, Adam C Krull, Vickie L Cooper, Terry J Engelken, Michael D Kleinhenz and Annette M OâConnor “Association between antimicrobial drug class for treatment and retreatment of bovine respiratory disease (BRD) and frequency of resistant BRD pathogen isolation from veterinary diagnostic laboratory samples” Number: 12 Publisher: Public Library of Science San Francisco, CA USA In PloS one 14.12, 2019, pp. e0219104
                                                                  • [13] Keith D. DeDonder and Michael D. Apley “A Review of the Expected Effects of Antimicrobials in Bovine Respiratory Disease Treatment and Control Using Outcomes from Published Randomized Clinical Trials with Negative Controls” Number: 1 In Veterinary Clinics of North America: Food Animal Practice 31.1, Bovine Clinical Pharmacology, 2015, pp. 97–111 DOI: 10.1016/j.cvfa.2014.11.003
                                                                  • [14] Alexis Delabouglise, Andrew James, Jean-FranÃğois Valarcher, Sara HagglÃijnd, Didier Raboisson and Jonathan Rushton “Linking disease epidemiology and livestock productivity: The case of bovine respiratory disease in France” Number: 12 In PLOS ONE 12.12, 2017, pp. e0189090 DOI: 10.1371/journal.pone.0189090
                                                                  • [15] T.A. Edwards “Control Methods for Bovine Respiratory Disease for Feedlot Cattle” In Veterinary Clinics of North America: Food Animal Practice 26.2, 2010, pp. 273–284 DOI: 10.1016/j.cvfa.2010.03.005
                                                                  • [16] P. Ezanno, M. Andraud, G. BeaunÃľe, T. Hoch, S. Krebs, A. Rault, S. Touzeau, E. Vergu and S. Widgren “How mechanistic modelling supports decision making for the control of enzootic infectious diseases” In Epidemics 32, 2020, pp. 100398 DOI: 10.1016/j.epidem.2020.100398
                                                                  • [17] G.. Frank, R.. Briggs and K.. Gillette “Colonization of the nasal passages of calves with Pasteurella haemolytica serotype 1 and regeneration of colonization after experimentally induced viral infection of the respiratory tract” Number: 8 In American Journal of Veterinary Research 47.8, 1986, pp. 1704–1707
                                                                  • [18] Robert W. Fulton, K. Blood, Roger J. Panciera, Mark E. Payton, Julia F. Ridpath, Anthony W. Confer, Jeremiah T. Saliki, Lurinda T. Burge, Ronald D. Welsh, Bill J. Johnson and Amy Reck “Lung Pathology and Infectious Agents in Fatal Feedlot Pneumonias and Relationship with Mortality, Disease Onset, and Treatments” Publisher: SAGE Publications Inc In Journal of Veterinary Diagnostic Investigation 21.4, 2009, pp. 464–477 DOI: 10.1177/104063870902100407
                                                                  • [19] Maria Gaudino, Brandy Nagamine, Mariette F. Ducatez and Gilles Meyer “Understanding the mechanisms of viral and bacterial coinfections in bovine respiratory disease: a comprehensive literature review of experimental evidence” Number: 1 In Veterinary Research 53.1, 2022, pp. 70 DOI: 10.1186/s13567-022-01086-1
                                                                  • [20] J.. GonzÃąlez-MartÃn, L. Elvira, M. CerviÃśo LÈ3pez, N. PÃľrez Villalobos, E. Calvo LÈ3pez-Guerrero and S. Astiz “Reducing antibiotic use: Selective metaphylaxis with florfenicol in commercial feedlots” In Livestock Science 141.2, 2011, pp. 173–181 DOI: 10.1016/j.livsci.2011.05.016
                                                                  • [21] Dee Griffin “Bovine Pasteurellosis and Other Bacterial Infections of the Respiratory Tract” Number: 1 In Veterinary Clinics of North America: Food Animal Practice 26.1, 2010, pp. 57–71 DOI: 10.1016/j.cvfa.2009.10.010
                                                                  • [22] Dee Griffin “The monster we don’t see: subclinical BRD in beef cattle” Number: 2 Publisher: Cambridge University Press In Animal health research reviews 15.2, 2014, pp. 138–141
                                                                  • [23] G.P. Grissett, B.J. White and R.L. Larson “Structured Literature Review of Responses of Cattle to Viral and Bacterial Pathogens Causing Bovine Respiratory Disease Complex” Number: 3 In Journal of Veterinary Internal Medicine 29.3, 2015, pp. 770–780 DOI: 10.1111/jvim.12597
                                                                  • [24] W. Hilton “BRD in 2014: where have we been, where are we now, and where do we want to go?” Number: 2 Publisher: Cambridge University Press In Animal Health Research Reviews 15.2, 2014, pp. 120–122 DOI: 10.1017/S1466252314000115